### v2

# CHAIR-Classifier of Hallucination As Improver

Abstract
In this work, we introduce CHAIR (Classifier of Hallucination As ImproveR), a supervised framework for detecting hallucinations by analyzing internal logits from each layer of every token. Our method extracts a compact set of features—such as maximum, minimum, mean, standard deviation, and slope—from the token logits across all layers, enabling effective hallucination detection without overfitting. Experiments on TruthfulQA and MMLU datasets demonstrate that CHAIR significantly improves detection accuracy, particularly in zero-shot scenarios, showcasing its robustness and generalizability.
Beyond hallucination detection, CHAIR highlights the potential of using internal representations for designing advanced decoding strategies. By leveraging patterns in logits, we suggest that more sophisticated models and adaptive decoding methods could further reduce hallucinations and enhance text completion quality. CHAIR not only offers a practical solution for detecting hallucinations but also lays the groundwork for exploring richer representations in LLMs to improve their factuality and coherence.
Index Terms:
Large language models, Representation learning
I Introduction
Hallucinations are a common issue in large language models (LLMs), limiting their reliability in critical applications, and a lot of research has been working on it [1] [2]. Hallucinations usually manifest as factual inaccuracies, misleading information, or even entirely fabricated content in model-generated text. For instance, when asked, ”What did Einstein invent?” the model might hallucinate and incorrectly associate Einstein with inventions like the ”lightbulb” or ”telephone,” which are actually Edison’s inventions. This issue not only affects the credibility of the model, but also poses potential risks in fields like healthcare and law.
Hallucinations can generally be divided into two types: Faithful and Truthful hallucinations. Faithful hallucinations usually involve external knowledge, focusing on whether the model can respond accurately to questions involving outside information. For example, the question ”What is the capital of France?” involves a faithful hallucination because it requires the model to rely on factual knowledge in its training data. Truthful hallucinations, on the other hand, refer to the model’s ability to produce accurate content based on its internal knowledge alone, even without external factual dependencies. For instance, answering a question like ”What kind of animal is a cat?” depends on the model’s internalized commonsense knowledge. This research  [^1] The source code for our research can be found at https://github.com/eggachecat/CHAIR. focuses mainly on truthful hallucinations, as they are directly related to the accuracy of the internal representations of the model.
I-A Related Work
Recent studies have proposed various innovative strategies to reduce hallucinations. Factual nucleus sampling[3] is one approach that constrains word selection based on probabilities during decoding, increasing the likelihood of generating factual correct content. For example, in response to ”How many satellites does Earth have?” limiting the vocabulary selection can guide the model to generate a response like ”one” or ”the Moon,” improving factual accuracy. Contrastive decoding [4] builds on this by contrasting different decoding paths to improve the reliability of the generated content. This approach focuses on tokens that initially receive low scores in the earlier layers but increase in score in later layers, suggesting that these tokens are syntactically plausible but semantically imprecise in the early stages, only becoming more accurate in later layers. For example, the model might initially lean toward generating ”Mars has plenty of water,” but later layers refine this to ”Mars has limited water resources.”
Another technique, DoLA[5], examines how different attention-layers contribute to hallucination. By analyzing attention scores across multiple layers, DoLA highlights which layers to be ”contrasted” tend to produce more factual and reliable outputs. This method shows that hallucination can sometimes arise from attention imbalances across layers and suggests that integrating attention information from multiple layers, rather than just relying on the final layer, can improve model accuracy.
Inference-Time Intervention [6] also addresses hallucinations by adjusting the model’s internal representations through vector shifting. For example, when answering ”What elements make up carbon dioxide?” this method can help the model accurately generate ”carbon and oxygen” rather than introducing unrelated elements. This approach aims to find a suitable adjustment direction during inference, helping align the internal representations with factual correctness.
In recent research, there has been increasing interest in leveraging the internal representations of LLMs. For example, [7] discusses how these internal layers can provide more informative features for downstream tasks compared to solely relying on the final layers. This insight highlights the potential benefits of exploring and utilizing information from internal representations, which aligns well with the objectives of our work in improving hallucination detection.
I-B Motivation
Building on these approaches and analyzing the model’s internal behavior on TruthfulQA [8] shown in Figure 1, we propose a new perspective: intermediate layer information is not only useful for syntactic structure but also provides additional signals for semantic accuracy. Solely relying on the final layer might be insufficient for fully capturing hallucinations, while integrating intermediate layer information could enhance response accuracy.
Motivated by this insight, we conducted the following research: for each token, we record its state information across every layer, forming a token-logit-trace. By analyzing these traces, we identified some notable patterns. For example, certain tokens might score lower in early layers but increase in later ones. This pattern suggests that while the model initially tends toward generating content related but incorrect, it progressively aligns with the accurate answer.
Using token-logit-traces, we performed supervised learning experiments on datasets such as TruthfulQA [8] and MMLU [9] designing a logistic regression classifier that aggregates token scores across all layers in LLaMA [10] or Mistral [11] into a compact set of key features. This classifier is then used to detect hallucinations in model outputs. Experimental results show that our method significantly improves performance in these datasets, especially in multiple-choice tasks, where accuracy is notably enhanced. This approach validates the effectiveness of intermediate layer information in hallucination detection and provides a strong foundation for future applications in identifying and mitigating hallucinations across various use cases.
Figure 1: This figure shows the score in different layers for each answer to two questions from TruthfulQA. The red lines represent scores trace, which is defined in Equation 3, of incorrect answers while the green lines indicate correct answers.In the left graph, although the green curves consistently have lower scores, their pattern of change is noticeably different from that of the red curves. This suggests that even when correct answers have lower scores, their variation across layers contrasts distinctly with that of the incorrect answers which is the information we could take advantage of. A similar phenomenon is observed in the right graph. Even though one of the green curves scores above most of the red ones, its shape closely resembles the other green curves, showing a difference from the red curves. This further demonstrates that correct and incorrect answers exhibit distinctly different score patterns across layers.
II Method
Our approach, named Classifier of Hallucination As Improver (CHAIR), leverages representations generated by each layer of a model structure similar to LLaMA to identify hallucinations in language model outputs. LLaMA is a transformer-based model that applies self-attention [12] and feed-forward layers iteratively to build contextual representations of tokens. A crucial component in this setup is the use of an lm_head layer, which maps the hidden state from the last layer to a score, representing the model’s confidence in each token.
To capture the sequential layer-wise representations, we define a function called history that maps each token to its hidden states across all layers. Additionally, we use the lm_head to compute a score for each layer’s output. This structure allows us to inspect both the hidden representations and the model’s token scores for each layer.
We now introduce the architecture of CHAIR first, then we will dive each component of this architecture.
II-A Architecture of CHAIR
Before we delve into the CHAIR architecture, let’s first revisit the structure of a LLaMA-like model. At its core, input tokens are converted into embeddings and passed through multiple layers of transformer blocks. Each block consists of layer normalization, a multi-head attention mechanism, and a feed-forward layer that processes each token independently. Layer normalization is applied before both the multi-head attention and feed-forward layers. Residual connections are incorporated to ensure stable training. The output from the final transformer layer is passed to a projection layer, referred to as the $lm\_head$.
Figure 2 illustrates the CHAIR model architecture, which is designed to handle multiple tokens at the feature extraction level. At the bottom, the input consists of an ”input embedding” comprising a question represented by a light gray rectangular box and multiple ”choice/answer tokens”. Each choice token is highlighted in a different color (e.g., blue and green) to signify distinct tokens.
The model contains multiple transformer blocks, identical to those in the LLaMA architecture, depicted as gray boxes labeled ”block-1” to ”block-N.” The output of each block is passed to an $lm\_head$ module, which applies a transformation specific to the corresponding block’s output.
After the $lm\_head$ transformation, the resulting outputs are treated as features for the corresponding token at that layer or block. Each colored feature corresponds to a specific token.
These token-specific features are aggregated and grouped, as indicated by dashed boxes. This grouping signifies that the features from all layers for each token are combined to construct a comprehensive feature set for that token.
Finally, the aggregated token-specific features are passed to a classifier. The classifier processes these features to determine whether the generated answer is hallucinatory.
Figure 2: This illustrates the structure of CHAIR. The blue and green represent the tokens for choices/answers, while the gray boxes denote the transformer layers in the LLaMA model. The purple rectangles correspond to the references of the $lm\_head$ layer, and the yellow rectangle represents the feature extraction module. The classifier processes the inputs to determine whether the output sequence represents a hallucination.
II-B LLaMA-like Model Structure
The LLaMA model consists of a series of Transformer layers, each containing a self-attention mechanism and a feed-forward neural network. Given a token at $t$-index representation at layer $l$, denoted as $h_{l}(t)$, the transformations in each layer are as follows:
1. Self-Attention Mechanism: The token representation is transformed to queries $Q$, keys $K$, and values $V$ using learned weights:
$Q$ $=h_{l-1}(t)W_{Q},$
$K$ $=h_{l-1}(t)W_{K},$
$V$ $=h_{l-1}(t)W_{V}$
where $W_{Q}$, $W_{K}$, and $W_{V}$ are weight matrices. The self-attention output is then:
$\text{ATT}(Q,K,V)=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$
where $d_{k}$ is the dimension of the queries.
2. Feed-Forward Network (FFN): The attention output is then passed through a feed-forward neural network:
$\text{FFN}(x)=\text{ReLU}(xW_{1}+b_{1})W_{2}+b_{2}$
3. Layer Output: The output at layer $l$, denoted $h_{l}(t)$, is computed using residual connections and layer normalization:
$h_{l}(t)=$ $\ \text{LN}\left(\text{ATT}(Q,K,V)+h_{l-1}(t)\right)$
$+\text{LN}\left(\text{FFN}(\text{ATT}(Q,K,V)+h_{l-1}(t))\right)$
II-C LM Head and History of Token’s Score
We define a function called $H$ (or $history$ ), defined in (1), that captures the sequential layer-wise representations of each token within the input sequence. This enables a detailed inspection of the model’s internal state as it processes each token, facilitating the identification of potential hallucinations based on abnormal layer-wise patterns.
$H:t\mapsto\{h_{1}(t),h_{2}(t),\dots,h_{L}(t)\}$ (1)
where $L$ is the layer count of the model.
In a LLaMA-like model, an lm_head layer is applied to the hidden state $h_{l}(t)$ at each layer to obtain a score for each token. This score, denoted $s(t)$, represents the logit associated with a particular token at position $t$:
$s_{i}(t)=\text{lm\_head}(h_{i}(t))$ (2)
And we define the Score function for a token at index $t$ with its $history$
$S:t\mapsto\{s_{1}(t),s_{2}(t),\dots,s_{L}(t)\}$ (3)
II-D Feature Extraction
To determine if a token at position $t$ is likely to be a hallucination, we analyze its H and $s$ through manually engineered features:
1. Final Logit (Last Layer Score):
$\text{Last Score}:=s_{L}(t)=\text{lm\_head}(h_{L}(t))$
2. Mean:
$\text{Mean}:=\frac{1}{L}\sum_{i=1}^{L}s_{i}(t)$
3. Maximum Value:
$\text{Max}:=\max(s_{1}(t),s_{2}(t),\dots,s_{L}(t))$
4. Minimum Value:
$\text{Min}:=\min(s_{1}(t),s_{2}(t),\dots,s_{L}(t))$
5. Standard Deviation:
$\text{Std}:=\sqrt{\frac{1}{L}\sum_{i=1}^{L}(s_{i}(t)-\text{Mean})^{2}}$
6. Slope:
$\text{Slope}:=\frac{\sum_{i=1}^{L}(i-\bar{i})(s_{i}(t)-\text{Mean})}{\sum_{i=1}^{L}(i-\bar{i})^{2}}$
The extracted features are then normalized. Let
[…]
The normalization is performed as follows:
$\text{features}_{\text{norm}}=\frac{\max(\text{features})-\text{features}}{\max(\text{features})-\min(\text{features})}$ (5)
This normalized feature vector, $\text{features}_{\text{norm}}$, serves as the input for a classifier $C$, mapping the features to a log-probability score that indicates the likelihood of hallucination.
II-E Supervised Learning
In this section, we describe the process of training the classifier. Constructing the training and testing datasets is relatively straightforward: for a given question, the expected answers are labeled as 1, while all other answers are labeled as 0.
For instance, as shown in Figure 1, the green answers are labeled as 1, whereas the red answers are labeled as 0.
The challenging part lies in how to sample from the dataset. We explored several approaches:
1. Dataset-based sampling: Following the method in [6], we sampled directly from the dataset used for testing. This approach resembles an n-shot learning paradigm: we randomly sample $n$ examples to create a training set, train the classifier on this set, and then evaluate its performance on the test set. This method is compared against other n-shot approaches.
2. Impact of $n$: To determine the effect of the number of samples $n$ on classifier performance, we conducted multiple experiments using the same $n$. The results are analyzed to identify the minimum $n$ required for robust classifier performance.
3. Cross-dataset generalization: We hypothesize that the logits from internal layers of the LLM reflect intrinsic characteristics of the model, which should generalize across datasets. Under this assumption, the logits themselves can reveal whether an output is hallucinatory. Once the classifier learns this pattern from one dataset, it should perform well on others, validating the generalizability of the approach.
These experiments collectively evaluate the effectiveness and robustness of the classifier in detecting hallucinations in LLM outputs.
III Experiment and Analysis
To comprehensively evaluate the effectiveness and robustness of our CHAIR method, we conducted a series of experiments addressing key research questions. First, we examined the model’s ability to generalize within the same dataset by using separate training and testing sets. This step ensures that the model effectively captures the underlying patterns necessary for hallucination detection without overfitting to the training data. Next, we assessed the method’s robustness on individual datasets, focusing on its stability under different training size. This evaluation demonstrates the model’s robustness. Finally, to explore the method’s generalizability, we performed cross-dataset experiments by training the model on one dataset and testing it on another. This analysis is critical for understanding the CHAIR method’s adaptability to diverse data distributions and its potential for real-world applications. The following subsections detail the results and insights from each of these experiments.
III-A Supervised Learning on Dataset
Table I presents a comparison of the performance of three approaches-baseline(Meta-Llama-3-8B-Instruct), DOLA, and our proposed method-across multiple metrics derived from the TruthfulQA and MMLU dataset. Each row corresponds to a specific metric, with columns showing the performance scores for each approach. Our method outperforms both Llama-3 and DOLA on the MC1 and MC3 metrics on TruthfulQA task, demonstrating a improvement in these areas. However, on the MC2 metrics, DOLA achieves relatively higher scores, though both DOLA and our method remain substantially better than the Llama-3.
TABLE I: Comparative Performance of Mistral, DoLA, and Our Method across TruthfulQA and MMLU Metrics.
BaseModel Task Metrics Baseline +DoLA +Ours
Llama-3-8B-Instruct TruthfulQA MC1 0.41 0.38 0.51
MC2 0.59 0.68 0.65
MC3 0.32 0.35 0.37
MMLU@0-shot Acc 0.45 0.45 0.48
MMLU@1-shot Acc 0.65 0.65 0.65
Mistral-7B-Instruct-v0.3 TruthfulQA MC1 0.48 0.38 0.53
MC2 0.66 0.65 0.68
MC3 0.37 0.35 0.39
MMLU@0-shot Acc 0.59 0.59 0.60
MMLU@1-shot Acc 0.60 0.60 0.60
III-B Robustness on single dataset
Figure 3 illustrates the relationship between training set size and the stability of our model’s performance, showing the amount of data required to reliably fit the model without overfitting or cherry-picking specific samples. To ensure robustness in the evaluation, we conducted multiple experiments for each specified data size. For example, for the smallest sample size of 1, we performed 50 trials-each time randomly selecting one data point as the training sample and using the remaining data for prediction.
The experiments were carried out for the following training set sizes: 1, 3, 5, 10, 15 and 20. Our findings show that after a threshold of 15 training samples, the occurrence of bad cases (i.e., poor predictions) becomes rare, suggesting that 15 training samples are sufficient to achieve stable model performance with high probability.
Furthermore, the results demonstrate the robustness of our approach, as the model maintains stability across different training set sizes. The decline in variance with increasing data size underscores the effectiveness of the method in utilizing small yet sufficient data for training without compromising reliability. This behavior validates the method’s potential for practical deployment in scenarios with limited data availability.
Figure 3: This illustrates the Impact of Training Set Size on Model Stability and Robustness, showing the relationship between the size of training data and the results, given the very limited number of parameters used in our CHAIR model. This plot shows the outcomes of 50 experiments for each sample size, where each trial randomly selects $k$ examples for training. The y-axis represents the improvement on MC1, while the x-axis shows the number of training examples. Each bar and scatter point represents the variation in performance improvement across the 50 trials for each sample size. This approach helps us understand how the amount of training data influences the model’s improvement on MC1 under limited parameter conditions.
III-C Generalization and Robustness cross multiple dataset
The Section III-C and Section III-C present the evaluation scores obtained when models are trained on one dataset (represented by rows) and evaluated on another dataset (represented by columns). Each cell $(i,j)$ indicates the change in performance when the model is trained on the dataset corresponding to row $i$ and tested on the dataset corresponding to column $j$. For example, the first column contains the change in MC1 scores, reflecting the variation in performance on TruthfulQA. In contrast, the second and third columns show the changes in average accuracy for the other datasets.
This table demonstrates how the model generalizes across datasets, highlighting performance shifts in both in-domain and out-of-domain scenarios compared to the Llama-3-8B-Instruct model. We observe that the model performs well on the MMLU and TruthfulQA datasets, achieving significant improvements. However, the results on the MMLU dataset appear less promising. We will further analyze the reasons behind this discrepancy to better understand and address the underlying issues.
TABLE II: Performance Gain over Meta-Llama-3-8B-Instruct Matrix: Training on Rows, Evaluating on Columns
Train Dataset Test Dataset
TruthfulQA
MMLU
ACC@0.45
TruthfulQA +0.11 +0.02
MMLU +0.09 +0.11
TABLE III: Performance Gain over Mistral-7B-Instruct-v0.3 Matrix: Training on Rows, Evaluating on Columns
Train Dataset Test Dataset
TruthfulQA
MMLU
ACC@0.59
TruthfulQA +0.01 +0.00
MMLU +0.05 +0.06
IV Conclusion
In this work, we presented CHAIR, a framework for detecting hallucinations in large language models (LLMs) by leveraging internal logits from each layer of every token. By using a lightweight supervised learning classifier, CHAIR demonstrated significant improvements in hallucination detection on datasets like TruthfulQA and MMLU, particularly excelling in zero-shot scenarios.
While the current approach has proven effective, there remains considerable potential for further advancements. The internal logits extracted from each layer offer rich, hierarchical representations that could be better exploited with more sophisticated models, such as transformer-based classifiers. These models might capture more complex interactions between layers and tokens, potentially improving classification accuracy and robustness across diverse datasets.
Moreover, our findings suggest that understanding the internal dynamics of LLMs could pave the way for developing better decoding strategies. By incorporating insights from logits patterns into decoding algorithms, we can guide the model toward generating more factual and coherent completions, reducing the likelihood of hallucinations during inference. Techniques such as adaptive decoding or contrastive reranking could integrate seamlessly with CHAIR to further enhance its utility in real-world applications.
By extending CHAIR with advanced models and decoding techniques, we can continue to improve the reliability and trustworthiness of language model outputs, enabling their broader adoption in critical domains.